Circular No. 7738

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 T4 (NEAT) S. Pravdo, E. F. Helin, M. Hicks, and K. Lawrence, Jet Propulsion

Laboratory, report the discovery by the NEAT program of a comet with a diffuse coma of diameter ∼ 4′′ and a southward tail ∼ 10′′ long on CCD images taken on Oct. 15 with the Palomar 1.2-m Schmidt telescope. Additional NEAT images on Oct. 21.4 UT show the comet as very diffuse and faint, elongated east-west.

2001 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 15.34693  1^{h}21^{m}22.^{s}08  −5^{o}18′27′′.8  20.3

All of the available astrometry, including apparent prediscovery observations made by LONEOS on Aug. 27, and the following orbital elements (from 17 observations, Aug. 27–Oct. 23) appear on MPEC 2001-U41.

T = 2002 Jan. 4.7457 TT  ω = 315.^{o}2964
e = 0.394952  Ω = 64.8321 }2000.0
q = 8.555151 AU  i = 15.2763

a = 14.139618 AU n^{o} = 0.0185373 P = 53.169 years

POSSIBLE NOVA IN THE SMALL MAGELLANIC CLOUD W. Liller, Viña del Mar, Chile, reports his discovery of a novalike object

(mag 12.1) on Tech Pan photographs taken on Oct. 21.0857 and 21.0879 UT. The new object, located at α = 0^{h}46^{m}27^{s}.8 ± 0^{s}.7, δ = −73^{o}29′40′′.1 ± 1′′.2 (equinox 2000.0), does not appear on any of 45 photographs taken by Liller since Sept. 2000, including 2001 Sept. 27.36 (limiting mag 13.5). This position is ∼ 1′.6 west and 1′.5 south of that of the Harvard variable HV1429 (= V355 SMC; Gen. Cat. Var. Stars, 5, 81). A CCD observation made by Liller on Oct. 22.0725 gave a broadband V magnitude of 12.22 (comparison star HD 4558). A. Pearce, Nedlands, W. Australia, reports that Liller’s possible nova appeared at visual mag 13.2 on Oct. 23.842.

NOVA IN M31 A. V. Filippenko and R. Chornock, University of California at Berkeley,

report that a CCD spectrum (range 330–1000 nm) obtained on Oct. 20 UT with the 3-m Shane reflector at Lick Observatory indicates that the apparent nova in M31 reported on IAUC 7729 is indeed a nova, with strong Balmer and Fe II emission lines.

2001 October 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green